LSAT test scores are normally distributed with a mean of 160 and a standard deviation of 7. What score would place you in the top 2% of test-takers? HINT [See Example 3.] (Round your answer to the nearest whole number.)
Answer #1

Solution :

Given that ,

mean = = 160

standard deviation = = 7

The z - distribution of the 2% is,

P( Z > z ) = 2%

1 - P( Z < z ) = 0.02

P( Z <  ) = 1 - 0.02

P( Z < z ) = 0.98

P( Z < 2.054 ) = 0.98

z = 2.054

Using z - score formula,

X = z * +

= 2.054 * 7 + 160

= 174.37

= 174